Which expression represent f(g(3)). If f(x)=3-x^2 & g(x)=1+2x?
A.25
B.1
C.-46
D.-11
Plz Help..

Answer by jim_thompson5910(35256)

f(x)=3-x^2

f(g(x))=3-(g(x))^2

f(g(x))=3-(1+2x)^2

f(g(3))=3-(1+2(3))^2

f(g(3))=3-(1+6)^2

f(g(3))=3-(7)^2

f(g(3))=3-49

f(g(3)) = -46